OpenStudy (anonymous):

If a won 80% of the games and they play 5 more games and won 3 and lost 2. Overall loss becomes 15%. Find out the total games played?

OpenStudy (aravindg):

Let total number of games played be x. (80/100)x =games won--->overall loss=(20/100)x Then they play 5 more games.So number of games played becomes x+5. Now 15/100(x+5)= New Overall loss

OpenStudy (anonymous):

thanks

OpenStudy (aravindg):

Did you get the final answer?

OpenStudy (aravindg):

Neglect my last step , we have to consider it with initial x and not x+5 (15/100)x=new overall loss (20/100)x+2=(15/100)x Solve this and you will get the answer.